OpenStudy (anonymous):

Estimate by rounding to the nearest hundred. 923 - 466

OpenStudy (asadkarim7):

900 - 500 = 400

OpenStudy (anonymous):

do they want you to round at the end or right away? you get different answers both ways

OpenStudy (anonymous):

oh never mind they are asking you to estimate, asadkarim is right

OpenStudy (saifoo.khan):

900 - 500 = 400

OpenStudy (anonymous):

Asad is right... the calculation does not become easy if we first find the difference and then round off...

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Estimate the answer by clustering. 725 + 676 + 712 + 729

OpenStudy (anonymous):

first two 700+700 = 1400 last two 700 + 700 = 1400 total 2800